Find the values given that loga 25 = 0.846 and loga 4 = 0.173:
Loga 20a _______________

Answer by stanbon(75887)

Find the values given that loga 25 = 0.846 and loga 4 = 0.173:
Loga 20a _______________
----
loga(20a) = loga(4) + loga(5) + loga(a)
------
Since loga(25) = 2loga(5) = 0.846, loga(5) = 0.423
And loga(a) = 1
------
Therefore, Loga(20a) = 0.173 + 0.423 + 1 = 1.596
-------------------------------------------------
Cheers,
Stan H.
==================